# Magneto-Thomson and transverse Thomson effects in an interacting hadron gas in the presence of an external magnetic field

### v1

@@ III Results and Discussions
The upper panel of Fig. 2 represents $Th_{B}$ as a function of $T$ for three different values of $\mu_{B}$ at 0.10, 0.30, and 0.40 GeV at magnetic field 0.1 $m_{\pi}^{2}$, whereas the lower panel represents those for the magnetic field 1.0 $m_{\pi}^{2}$. For the case of $eB$ = 0.1 $m_{\pi}^{2}$, we observe that $Th_{B}$ decreases with the temperature up to 0.1 GeV, becomes zero at a particular $T$, and further, it starts to increase in the negative direction for higher values of temperature. Similar to what was observed in the case of a vanishing magnetic field, all the models agree up to a certain temperature and then start deviating from each other at higher temperatures. Furthermore, we also observe that the values of $Th_{B}$ increase in the negative direction with the increasing values of $\mu_{B}$ for higher $T$ but decrease in the positive direction for the lower values of $T$. Unlike the case without a magnetic field, the magneto-Thomson coefficient $Th_{B}$ exhibits an additional dependence on the Hall-like component of electrical conductivity and other integral terms that arise due to the presence of an external magnetic field. The Hall-like conductivity term, $\sigma_{H}/T$, increases with baryon chemical potential $\mu_{B}$ because of the growing net baryonic contributions. It is important to note that mesons do not contribute directly to the integrals $\mathcal{I}_{31}/T^{2}$ and $\mathcal{I}_{42}/T^{2}$, though their influence appears indirectly through the enthalpy per baryon $h$. The effect of the magnetic field is significant on $\mathcal{I}_{31}/T^{2}$ at low temperatures, but almost diminishes at higher temperatures due to a reduced relaxation time. On the other hand, $\mathcal{I}_{42}/T^{2}$ first increases with temperature and then starts to decrease. The higher the magnetic field, the lower the $\mathcal{I}_{42}/T^{2}$. Here, the magnitude of $\sigma_{el}/T$ is significantly larger than that of $\sigma_{H}/T$, the integrals $\mathcal{I}_{31}/T^{2}$ and $\mathcal{I}_{42}/T^{2}$ are of comparable scale [63]. Consequently, the magneto-Seebeck coefficient $S_{B}$ can be approximately represented by the ratio $(\mathcal{I}_{31}/T^{2})/(\sigma_{el}/T)$, and its temperature dependence is reflected in the behavior of the magneto-Thomson coefficient $Th_{B}$, as shown in Fig. 2. The $T$ and $\mu_{B}$ dependence of $\sigma_{el}/T$ in a magnetic field causes $S_{B}$ to increase at low $T$ and decrease at high $T$. This results in a positive $Th_{B}$ at lower temperatures, which decreases with $T$ and becomes negative at higher temperatures. At low $T$, a large value of $\Omega_{c_{i}}\tau_{R_{i}}$ suppresses the mesonic contribution to $\sigma_{el}/T$, and this reduction is compensated by the enhanced baryonic contribution as $\mu_{B}$ increases. However, at higher temperatures, the baryonic increase is insufficient to compensate for the mesonic suppression, leading to an overall decrease in $\sigma_{el}/T$ with $\mu_{B}$. For stronger magnetic fields, as shown in the lower panel of Fig. 2, $Th_{B}$ shows large and positive values at low $T$, gradually approaching zero at high $T$. It is observed that for higher values of a magnetic field, results from all the models agree up to $T\approx 0.15$ GeV at $\mu_{B}$ = 0.10 GeV. With increasing values of $\mu_{B}$, the disagreement between the models starts from even lower values of $T$. This observation is in line with what was observed for the magneto-Seebeck coefficient in Ref. [13].
[…]
@@ IV Summary
In heavy-ion collisions, temperature gradients between the central and peripheral regions of the created fireball, especially in the presence of a non-zero baryon chemical potential, can introduce thermoelectric effects within the hot hadronic medium. The presence of a magnetic field induces anisotropy in the system by breaking rotational symmetry, enhancing direction-dependent thermoelectric responses. In this work, for the first time, we estimate the magneto-Thomson and transverse Thomson coefficients using different HRG models in the presence of a magnetic field. These higher-order thermoelectric coefficients arise from the temperature dependence of the leading-order thermoelectric coefficients. The Thomson coefficient quantifies the net generation or absorption of heat in the medium when current flows in the medium in the presence of a temperature gradient. We have studied this phenomenon in the presence of an external magnetic field. Here, the baryon conservation is considered for the heat conduction in the hot hadronic medium. Using the Gibbs-Duhem relation, we relate the gradients in baryon chemical potential to those in temperature, revealing the role of enthalpy in modifying the thermoelectric coefficients. The net leading-order thermoelectric coefficient vanishes for mesons due to their charge symmetry; mesons still have a significant effect on the leading-order thermoelectric coefficients because of their contributions to enthalpy. Furthermore, we explore the effect of different values of external magnetic fields on these transport properties. It is observed that in the presence of a higher magnetic field ($eB=1.0~{}m_{\pi}^{2}$), both the $Th_{B}$ and $Th_{N}$ show higher values as compared to lower values of magnetic field ($eB=0.1~{}m_{\pi}^{2}$). It is also observed that for both the coefficients, the effect of different models is mostly significant at higher temperatures. The results for all four models almost overlap at the lower temperatures. In the absence of a magnetic field, the magneto-Thomson coefficient $Th_{B}$ reduces to the Thomson coefficient $Th$. On the other hand, the transverse Thomson coefficient $Th_{N}$ completely vanishes in the absence of a magnetic field because of the vanishing Lorentz force. Recently, both phenomena have also been studied in Ref. [62] for the QGP medium in detail using the quasi-particle approach.

### v2

@@ III Results and Discussions
The upper panel of Fig. 2 represents $Th_{B}$ as a function of $T$ for three different values of $\mu_{B}$ at 0.10, 0.30, and 0.40 GeV at magnetic field 0.1 $m_{\pi}^{2}$, whereas the lower panel represents those for the magnetic field 1.0 $m_{\pi}^{2}$. For the case of $eB$ = 0.1 $m_{\pi}^{2}$, we observe that $Th_{B}$ decreases with the temperature up to 0.1 GeV, becomes zero at a particular $T$, and further, it starts to increase in the negative direction for higher values of temperature. Similar to what was observed in the case of a vanishing magnetic field, all the models agree up to a certain temperature and then start deviating from each other at higher temperatures. Furthermore, we also observe that the values of $Th_{B}$ increase in the negative direction with the increasing values of $\mu_{B}$ for higher $T$ but decrease in the positive direction for the lower values of $T$. Unlike the case without a magnetic field, the magneto-Thomson coefficient $Th_{B}$ exhibits an additional dependence on the Hall-like component of electrical conductivity and other integral terms that arise due to the presence of an external magnetic field. The Hall-like conductivity term, $\sigma_{H}/T$, increases with baryon chemical potential $\mu_{B}$ because of the growing net baryonic contributions. It is important to note that mesons do not contribute directly to the integrals $\mathcal{I}_{31}/T^{2}$ and $\mathcal{I}_{42}/T^{2}$, though their influence appears indirectly through the enthalpy per baryon $h$. The effect of the magnetic field is significant on $\mathcal{I}_{31}/T^{2}$ at low temperatures, but almost diminishes at higher temperatures due to a reduced relaxation time. On the other hand, $\mathcal{I}_{42}/T^{2}$ first increases with temperature and then starts to decrease. The higher the magnetic field, the lower the $\mathcal{I}_{42}/T^{2}$. Here, the magnitude of $\sigma_{el}/T$ is significantly larger than that of $\sigma_{H}/T$, the integrals $\mathcal{I}_{31}/T^{2}$ and $\mathcal{I}_{42}/T^{2}$ are of comparable scale [70]. Consequently, the magneto-Seebeck coefficient $S_{B}$ can be approximately represented by the ratio $(\mathcal{I}_{31}/T^{2})/(\sigma_{el}/T)$, and its temperature dependence is reflected in the behavior of the magneto-Thomson coefficient $Th_{B}$, as shown in Fig. 2. The $T$ and $\mu_{B}$ dependence of $\sigma_{el}/T$ in a magnetic field causes $S_{B}$ to increase at low $T$ and decrease at high $T$. This results in a positive $Th_{B}$ at lower temperatures, which decreases with $T$ and becomes negative at higher temperatures. At low $T$, a large value of $\Omega_{c_{i}}\tau_{R_{i}}$ suppresses the mesonic contribution to $\sigma_{el}/T$, and this reduction is compensated by the enhanced baryonic contribution as $\mu_{B}$ increases. However, at higher temperatures, the baryonic increase is insufficient to compensate for the mesonic suppression, leading to an overall decrease in $\sigma_{el}/T$ with $\mu_{B}$. For stronger magnetic fields, as shown in the lower panel of Fig. 2, $Th_{B}$ shows large and positive values at low $T$, gradually approaching zero at high $T$. It is observed that for higher values of a magnetic field, results from all the models agree up to $T\approx 0.12$ GeV at $\mu_{B}$ = 0.10 GeV. With increasing values of $\mu_{B}$, the disagreement between the models starts from even lower values of $T$. This observation is in line with what was observed for the magneto-Seebeck coefficient in Ref. [13].
Figure 4: Upper panel: left figure shows scaled electrical conductivity ($\sigma_{el}/T$) and right figure shows scaled Hall-like component of electrical conductivity ($\sigma_{H}/T$), bottom panel: left figure shows scaled integral ($\mathcal{I}_{31}/T^{2}$) and right figure shows scaled integral ($\mathcal{I}_{42}/T^{2}$) at $\mu_{B}$ = 0.30 GeV with three different scenario of magnetic field eBeB = 0.0, eBeB = 0.1 $m_{\pi}^{2}$, and $eB_{0}$ = 1.0 $m_{\pi}^{2}$ with decay parameter $\tau_{B}$ = 6 fm.
Figure 5: Left panel: Magneto-Thomson coefficient ($Th_{B}$), and right panel: transverse Thomson coefficient ($Th_{N}$) as a function of temperature for baryon chemical potential at $\mu_{B}$ = 0.30 GeV with three different scenario of magnetic field eBeB = 0.0, eBeB = 0.1 $m_{\pi}^{2}$, and $eB_{0}$ = 1.0 $m_{\pi}^{2}$ with decay parameter $\tau_{B}$ = 6 fm.
In the presence of an external magnetic field, we also observe a Hall-like component for the higher-order thermoelectric coefficients. The presence of a non-zero temperature gradient in a conducting medium directed perpendicular to the direction of a magnetic field gives rise to the Nernst coefficient. It measures the electric current perpendicular to both the magnetic field and the temperature gradient. In the absence of a magnetic field, this coefficient vanishes. The temperature dependence of the Nernst coefficient in the medium further gives rise to the transverse Thomson coefficient $Th_{N}$ as shown in Eq. (77). Fig. 3 represents $Th_{N}$ as a function of $T$ for three different values of $\mu_{B}$ at 0.10, 0.30, and 0.40 GeV. The upper and lower panels present the results for the case of the magnetic field with values 0.1 and 1.0 $m_{\pi}^{2}$, respectively. In the upper panel, we observe that $Th_{N}$ is positive for lower values of $T$ up to 0.10 GeV for $\mu_{B}$ = 0.10 and 0.30 GeV. The positive values of $Th_{N}$ signify that in Eq. (77), the quantity 2NBNB is dominating over the dynamical term, $T\frac{dNB}{dT}$ at lower temperatures. The impact of this dynamic part slowly starts increasing with $T$, and hence the negative behavior of $Th_{N}$ is also observed at higher $T$. For the higher values of $\mu_{B}$, it is observed that the dynamical part of Eq. (77) is dominating almost for the whole temperature range considered here. As it is observed in Ref. [13], the coefficient $NB$ vanishes near high temperatures; here, $Th_{N}$ also starts disappearing at higher temperatures. For $eB$ = 1.0 $m_{\pi}^{2}$, results from different models are observed to be almost similar, ı.e., no significant model dependence is observed. Even at higher $\mu_{B}$, only a slight deviation in the models can be seen. As one goes to a higher value of the magnetic field ($eB$ = 1.0 $m_{\pi}^{2}$), $Th_{N}$ is observed to be higher. Due to the higher order of magnitude of $\sigma_{el}/T$ as compare to $\sigma_{H}/T$, $NB$ can be approximated by ratio $(\mathcal{I}_{42}/T^{2})/(\sigma_{el}/T)$. The sharp decrease of $\mathcal{I}_{42}/T^{2}$ as compare to $\sigma_{el}/T$ give rise to decreasing trend of $NB$ with $\mu_{B}$. Though $NB$ decreases with increasing $\mu_{B}$ but the dynamical rate of decrease of $NB$ increases. Hence, the coefficient $Th_{N}$ mostly has negative values throughout the whole temperature range at high values of $\mu_{B}$. Here, a clear effect of magnetic field values can be observed in the upper and bottom panels of Fig. 3. For the case of a higher magnetic field, the effect of the Lorentz force is also high. Thus, the higher Lorentz force increases the thermoelectric effects in the transverse direction, which results in higher values of the transverse Thomson coefficient. Once this magnetic field vanishes, the coefficient $Th_{N}$ drops to zero only.
For completeness, we further extend this study to the case of a time-varying electromagnetic field. Here, as mentioned in Eq. 86, we have considered the exponential profile of the electromagnetic field. As the hadronic model considered in this study represents a static medium, we observe the effects of a time-varying magnetic field in a snapshot at proper time $\tau$ = 9 fm. The results presented in Figs. 4 and 5 are focused on examining the influence of a time-varying magnetic field on the transport behavior of the hadronic medium. As the magnetic field decays and weakens over time, the effects of Landau quantization become negligible. Therefore, for a consistent comparison between the dynamical and static magnetic field pictures, Landau quantization has not been included in either case in these plots, allowing us to keep the focus on the dynamical effects of the temporal evolution of the field. As shown in Fig. 4, the upper-left panel displays the variation of the scaled electrical conductivity $\sigma_{el}/T$ with temperature, while the upper-right panel presents the Hall-like component $\sigma_{H}/T$. The lower panels correspond to the scaled heat current integrals $\mathcal{I}_{31}/T^{2}$ and $\mathcal{I}_{42}/T^{2}$, which constitute the underlying inputs for evaluating the leading-order and higher-order thermoelectric coefficients. Three distinct scenarios are compared: the case without a magnetic field ($eB=0.0$, black solid line), a constant magnetic field ($eB=1.0\,m_{\pi}^{2}$, cyan dashed line), and a time-dependent magnetic field with decay parameter $\tau_{B}=6~\mathrm{fm}$ (red dotted line). At zero magnetic field, $\sigma_{el}/T$ exhibits a monotonically decreasing trend with increasing temperature, consistent with the expected reduction of medium interaction effects at higher temperatures [16]. The introduction of a magnetic field modifies the charge transport behavior, leading to an overall suppression in $\sigma_{el}/T$ at low temperatures due to the cyclotron motion of charged hadrons. The damping effect for the case of a decaying magnetic field is smaller due to the reduction in the magnitude of the magnetic field. In contrast, the Hall-like component $\sigma_{H}/T$ shows a non-monotonic structure, attaining a peak in the low-temperature regime where the Lorentz force contribution is dominant. The peak shifts towards lower temperatures and decreases in magnitude when the field becomes time dependent, indicating a dynamical competition between the magnetic relaxation time and the microscopic relaxation processes.
The thermal integrals $\mathcal{I}_{31}/T^{2}$ and $\mathcal{I}_{42}/T^{2}$ exhibit complementary trends. The integral $\mathcal{I}_{31}/T^{2}$ remains negative throughout the temperature range, gradually approaching zero as the temperature increases, while $\mathcal{I}_{42}/T^{2}$ shows a pronounced peak structure analogous to that of $\sigma_{H}/T$. These behaviors underline the sensitivity of heat flow to both the strength and temporal variation of the magnetic field. Overall, the comparison among the three cases shows that the presence of a time-varying magnetic field reduces the magnitude of both electrical and thermal transport in the hadronic medium. The decaying field weakens the response of charged particles to the electromagnetic force, resulting in smoother and less pronounced variations of the transport coefficients with temperature. This suggests that time-dependent magnetic effects can significantly impact the charge and heat transport properties of the hadronic matter produced in heavy-ion collisions, providing a more realistic description of the medium in evolving magnetic backgrounds.
The results presented in Fig. 5 show the temperature dependence of the magneto-Thomson coefficient $Th_{B}$ and the transverse Thomson coefficient $Th_{N}$ for different magnetic field configurations in the hadronic medium at $\mu_{B}=0.30~\text{GeV}$. The integrals $H_{1_{i}}$$H_{4_{i}}$, shown in Eqs. 9396, encapsulate the effects of electric and heat current responses in the presence of both constant and time-dependent magnetic fields, and serve as the essential building blocks for evaluating magneto-Seebeck coefficient $S_{B}$ and normalized Nernst coefficient $NB$. The left panel shows that $Th_{B}$ decreases rapidly in magnitude with temperature for all cases, indicating a diminishing magneto-thermoelectric response in the high-temperature regime. At zero magnetic field, $Th_{B}$ remains relatively constant and small in magnitude, while the introduction of a finite magnetic field ($eB=1.0\,m_{\pi}^{2}$) enhances the response significantly at lower temperatures. When the magnetic field is allowed to decay in time with a decay parameter $\tau_{B}=6$ fm, the enhancement is notably suppressed, leading to a less pronounced temperature dependence. This demonstrates that the temporal damping of the magnetic field reduces the magnetically induced thermoelectric effects in the hadronic medium. The right panel shows $Th_{N}$, which reflects the Hall-like component of higher-order thermoelectric transport. Its temperature dependence arises from the interplay between the temperature derivative of the normalized Nernst coefficient and its magnitude. In the absence of a magnetic field, $Th_{N}$ remains zero across the entire temperature range due to the absence of $NB$. When a static magnetic field is introduced, $Th_{N}$ attains a negative value in the lower temperature region, indicating the dynamical part $T\frac{dNB}{dT}$ over $NB$ is dominant in Eq. 77. In the case of a time-varying magnetic field, the overall value of $Th_{N}$ is notably positive and decreases to zero at higher temperatures. Overall, these results reveal that the leading transport coefficients, such as $\sigma_{el}$ and $\sigma_{H}$, have relatively less sensitivity towards a time-varying magnetic field as compared to those of higher-order thermoelectric transport coefficients. This enhanced sensitivity arises from their dependence on temperature derivatives and mixed charge–heat transport terms, making them more responsive to the temporal evolution of the magnetic field.
@@ IV Summary
In heavy-ion collisions, temperature gradients between the central and peripheral regions of the created fireball, especially in the presence of a non-zero baryon chemical potential, can introduce thermoelectric effects within the hot hadronic medium. The presence of a magnetic field induces anisotropy in the system by breaking rotational symmetry, enhancing direction-dependent thermoelectric responses. In this work, for the first time, we estimate the magneto-Thomson and transverse Thomson coefficients using different HRG models in the presence of a magnetic field. These higher-order thermoelectric coefficients arise from the temperature dependence of the leading-order thermoelectric coefficients. The Thomson coefficient quantifies the net generation or absorption of heat in the medium when current flows in the medium in the presence of a temperature gradient. We have studied this phenomenon in the presence of an external magnetic field. Here, the baryon conservation is considered for the heat conduction in the hot hadronic medium. Using the Gibbs-Duhem relation, we relate the gradients in baryon chemical potential to those in temperature, revealing the role of enthalpy in modifying the thermoelectric coefficients. The net leading-order thermoelectric coefficient vanishes for mesons due to their charge symmetry; mesons still have a significant effect on the leading-order thermoelectric coefficients because of their contributions to enthalpy. Furthermore, we explore the effect of different values of external magnetic fields on these transport properties. It is observed that in the presence of a higher magnetic field ($eB=1.0~m_{\pi}^{2}$), both the $Th_{B}$ and $Th_{N}$ show higher values as compared to lower values of magnetic field ($eB=0.1~m_{\pi}^{2}$). It is also observed that for both the coefficients, the effect of different models is mostly significant at higher temperatures. The results for all four models almost overlap at the lower temperatures. In the absence of a magnetic field, the magneto-Thomson coefficient $Th_{B}$ reduces to the Thomson coefficient $Th$. On the other hand, the transverse Thomson coefficient $Th_{N}$ completely vanishes in the absence of a magnetic field because of the vanishing Lorentz force. Recently, both phenomena have also been studied in Ref. [65] for the QGP medium in detail using the quasi-particle approach. The study is extended to include a time-varying magnetic field with an exponential decay profile, allowing an observation of its impact on the dynamics of the hadronic medium. The electrical conductivity and related heat current integrals show that a decaying magnetic field leads to smoother and weaker transport responses compared to the constant magnetic field case. The magneto-Thomson and transverse Thomson coefficients further show that the higher-order thermoelectric transport is strongly influenced by magnetic field dynamics. Concisely, the higher-order thermoelectric coefficients exhibit a significantly stronger sensitivity to time-varying magnetic fields than the leading-order transport coefficients such as electrical conductivities and heat currents, highlighting the importance of magnetic field evolution in realistic heavy-ion collision dynamics.